OpenStudy (anonymous):

How do I solve : 3/10x - 1/5= -2

OpenStudy (anonymous):

add 1/5 to both sides

OpenStudy (anonymous):

and then multiply by the recipricol of 3/10

OpenStudy (anonymous):

is it (3/10)x - 1/5 = -2?

OpenStudy (anonymous):

x = 6?

OpenStudy (anonymous):

its easier to make the terms similar so make -1/5 -2/10

OpenStudy (anonymous):

does the answer has to be fraction?

OpenStudy (anonymous):

add 2/10 from both sides

OpenStudy (anonymous):

so now you have 3/10x=-2+2/10

OpenStudy (anonymous):

now multiply both sides by 10/3

OpenStudy (anonymous):

and you have your answer

OpenStudy (anonymous):

The answer should be -6. I just want to figure out an easy way to solve.

OpenStudy (anonymous):

@Trexy no but it should be -6. Could you show me in sketch? Sorry I'm a visual learner :)

OpenStudy (anonymous):

ok this how i Solve it but I dont know how people solve the other way. so I got